OpenStudy (unklerhaukus):

find the second derivative of y with respect to x C,D are constants \[y=(C+Dx)xe^{-2x}\]

OpenStudy (unklerhaukus):

what is the easy way to do this

OpenStudy (unklerhaukus):

\[y^\prime=(Dx+C+Dx-2x(C+Dx))e^{-2x}\]\[\qquad\qquad =(Dx+C+Dx-2Cx-2Dx)e^{-2x}\]

OpenStudy (anonymous):

This one method that you are solving is easier one.. If we distribute and then solve then it will becomes hard and lengthy too.. Always try to make factors when using Product Rule..